OpenStudy (anonymous):

The area of a room is x^2+9x+18 square feet. What is the length and width of the room? Express the answer in terms of x.

OpenStudy (anonymous):

(x+3)(x+6) length = x +6 width = x+ 3

OpenStudy (anonymous):

ok thanks :)